OpenStudy (anonymous):

How to factor: y= 2(x-3)^2 - 6

OpenStudy (anonymous):

u mean solve?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

is it y or 0?

OpenStudy (anonymous):

its y

OpenStudy (anonymous):

the only think i can think of is to use completing the square method to find the value of x and y when the curve cuts the y-axis

OpenStudy (anonymous):

what topic is this in math? if i know the topic it will help me

OpenStudy (anonymous):

Quadratic Functions

OpenStudy (anonymous):

you have to solve for x and y right? also r u familiar with completing the square method?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok what u do is state Let, x-3=0 therefore x=3 and y=2(0)^2-6 y=-6 understandable?

OpenStudy (anonymous):

ya thnx